OpenStudy (anonymous):

Write a two-column proof. Given: PR and QS bisect each other. Prove: Triangle PQR is congruent to Triangle RSP.

OpenStudy (anonymous):

Statements and Reasons?

OpenStudy (anonymous):



OpenStudy (anonymous):

Yes. Statements and Reasons. You know, a proof. I understand most of the proofs, but just not THIS one.

OpenStudy (anonymous):

Okay just a sec then :)

OpenStudy (anonymous):

ok. Thanks

OpenStudy (anonymous):

Did you take about parallelograms?

OpenStudy (anonymous):

I know it starts off like this.... Statements Reasons PR and QS bisect each other GIVEN PT is congruent to TR, SEGMENT BISECTOR QT is congruent to TS

OpenStudy (anonymous):

Oh I SOOOO hate this way lol / you didn't take about parallelograms right?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

so can you help me?

OpenStudy (anonymous):

I think so :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Hey, are you there?

OpenStudy (anonymous):

Yes soorrrryy!

OpenStudy (anonymous):

its k

OpenStudy (anonymous):

What's the name of the chapter you're in now?

OpenStudy (anonymous):

gimme a min

OpenStudy (anonymous):

Congruent Triangles

OpenStudy (anonymous):

HELLO???

OpenStudy (anonymous):

\[\cos^{SMALL wingspan} \int\limits_{YOUR AN}^{retrice \sqrt{flutter YOU feather}= YOU. UR. MUTHA. flutterER\]